# WikiDBGraph: A Data Management Benchmark Suite for Collaborative Learning over Database Silos thanks: *Equal Contribution

### v2

@@ I Introduction
The CL pipeline is illustrated in Fig. 1 using an example involving a bank and an e-commerce company. The bank maintains a database of customers’ names and transaction records, while the e-commerce company maintains users’ purchase histories. The two clients can collaborate to train a model that recommends items at prices aligned with users’ income and consumption habits. This is achieved without sharing private data by exchanging intermediate results such as model parameters, gradients, or embeddings.
@@
By investigating the recently released large-scale database corpus WikiDBs [16], we identify three common yet impractical assumptions in existing CL benchmarks and algorithm design: (1) isolation: each database is treated as an isolated data source, and connections between clients are ignored; (2) alignment: databases are assumed to be aligned either horizontally (disjoint samples, same features) [17] or vertically (disjoint features, same samples) [18]; and (3) joinable: databases can be fully joined into a single database for training.
@@
We address these challenges by constructing WikiDBGraph, a large-scale, open-source graph of relational databases built from Wikidata [19]. To mitigate the limited explicit correlations, we leverage known relations within WikiDBs [16] and employ contrastive learning to train a model that predicts similarity between databases. The resulting similarity scores induce a graph whose nodes represent databases and whose weighted edges quantify their correlation. To address unaligned and unjoinable databases, we enrich the graph by defining and computing a set of properties for both nodes and edges, derived from database structure, schemas, and data distributions. Using these properties, we propose an automated data-mining pipeline for CL. While this pipeline yields average performance improvements over training on isolated databases, it reveals a critical performance gap compared to ideal centralized training.
The contributions of this work include:
@@
    We construct WikiDBGraph, a large-scale graph of 100,000 interconnected relational databases, to serve as a real-world dataset for collaborative tabular learning.
@@
    We enrich the graph with a comprehensive set of 13 node and 12 edge properties derived from database structure, schema, and content.
@@
  4. 4.
    We conduct case studies on feature- and instance-overlapped database pairs, summarizing key challenges to guide future research.
@@
The paper is organized as follows. Section II surveys database corpora and CL datasets and positions WikiDBGraph; Section III details construction—problem formulation, a contrastive similarity model, and graph assembly; Section IV analyzes graph structure and node/edge properties; Section V evaluates utility via an automated data-mining pipeline; Section VI presents feature-, instance-, and hybrid-overlap CL case studies; Section VII discusses findings and open challenges; Section VIII presents the future directions for data management; Section IX concludes the paper.
II Related Work
@@
Collaborative Learning
To address the data-silo problem, federated learning (FL) [17] and specific frameworks are proposed to learn models without sharing raw data [5, 6, 8, 7]. Beyond FL, collaborative learning (CL) systems such as FATE [22] unify multiple privacy-preserving paradigms—including FL, split learning, and transfer learning—and support diverse models, including linear regression, gradient-boosting decision trees, and support vector machines. Despite this progress, real-world CL deployments remain limited, largely because most algorithms target the model-training stage rather than the end-to-end data management pipeline.
@@
Existing collaborative learning (CL) datasets, particularly for federated learning (FL) and split learning (SL), inadequately capture the complexity of real-world, database-centric collaboration. Many benchmarks are synthetic, created by partitioning a centralized dataset to simulate distribution, including FedNoisy [23], FPLLib [24], MarsFL [25], NIIDBench [4], SLPerf [26], and VertiBench [13]. These synthetic partitions may not reflect real data distributions. Other datasets are derived from real-world sources—such as LEAF [11], ORAF [12], FedML [14], and NUS-WIDE [27]—but assume perfect alignment into horizontal CL (disjoint samples, same features) or vertical CL (disjoint features, same samples) [17]. This assumption often fails when both data and features overlap only partially. Moreover, these benchmarks typically (i) represent each client by a single table and (ii) treat inter-client relationships as uniformly important. Both assumptions are restrictive: client databases may be unjoinable into a single table, and inter-client relationships can vary substantially. Built from real-world Wikidata, WikiDBGraph is designed to capture this complexity, exposing practical CL challenges across clients with hybrid relationships and varying degrees of correlation.
@@
WikiDBGraph Real Horizontal and Verical Database
@@
Existing corpora of tabular data have evolved from collections of individual tables, such as GitTables [30], whose focus on isolated tables overlooks the relational nature of real-world data. To better reflect practical scenarios, database corpora have emerged, featuring multiple interrelated tables that support complex relational tasks. These resources vary in composition: schema-only corpora like SchemaDB [31] and GitSchemas [32] offer detailed schemas with minimal or no data; data-only corpora such as SQLShare [29] provide table content, often without explicit schemas; and schema-data corpora contain both. Within this last category, many are either small-scale, like the CTU Prague Relational Learning Repository [34], or tailored for specific tasks like Text-to-SQL (e.g., Spider [35], BIRD [33]), which limits their utility for general-purpose representation learning. Even recent large-scale schema-data corpora such as SchemaPile [36] and WikiDBs [16] share a critical shortcoming: the scarcity of explicitly defined relationships between databases. This gap has been a major impediment to studying collaborative learning across distinct database systems. Our work, WikiDBGraph, addresses this limitation by systematically constructing inter-database connections, thereby enabling research into large-scale cross-database collaborative learning.
III Dataset Construction
This section details the methods to construct WikiDBGraph. We begin by formally defining the problem of database relationship identification (Section III-A). Subsequently, our proposed solution is presented in Section III-B. The efficacy of this approach in identifying meaningful correlations is then evaluated in Section III-C. Finally, Section III-D describes the utilization of the validated model to construct the database graph. The pipeline of WikiDBGraph construction is shown in Fig. 2.
III-A Problem Definition
@@ III-B Approach
To mitigate the verbosity of the original JSON files and reduce input size for subsequent processing, we serialize each $s_{i}$ and samples in $D_{i}$ into a concise textual format, denoted as abstract $t_{i}\in\mathcal{T}$. This serialization retains both structural information and an abstract of the data content. Specifically, for each database, we preserve its name, the names of its tables, and for each table, the names of its columns along with a few representative sample values from each column. An illustrative example of our serialization format is presented below:
@@
While sample values are included to offer a qualitative indication of the data, we do not incorporate full data distributions or the entirety of the raw data. This approach is adopted for two primary reasons: 1) schema-level information (database, table, and column names) combined with data samples is often more directly indicative of the database’s topic than comprehensive statistical distributions of all values. 2) Processing and representing the complete data for all databases would be computationally expensive and lead to excessively large serialized representations.
@@
We employ a contrastive learning [37] framework to train an embedding model, $f_{\theta}(\cdot):\mathcal{T}\rightarrow\mathbb{R}^{d}$. This model maps serialized database schema representations from the input space $\mathcal{T}$ into a $d$-dimensional vector space. The distinction between positive and negative pairs, crucial for the contrastive loss, is determined by TIDs: a pair of database abstracts $(t_{i},t_{j})$ is a positive pair if they share an identical TID; otherwise, it is a negative pair. The sets of positive and negative pairs are partitioned into training, validation, and test subsets in a 7:1:2 ratio, ensuring no database overlap between partitions.
@@
TABLE III: Performance of the embedding model (BGE-M3) on the test set with threshold 0.5.
Model Performance
AUC-ROC F1 Precision Recall
Original 0.9651±0.0005 0.5533±0.0028 0.3856±0.0027 0.9790±0.0000
Fine-tuned 0.9987±0.0001 0.9270±0.0023 0.8687±0.0041 0.9938±0.0000
@@
III-C Evaluation of Embedding Model
[…]
@@ III-C Evaluation of Embedding Model
In addition, when compared to explicit positive pairs (those sharing a TID), this analysis reveals a substantial number of previously unidentified related databases exhibiting high similarity scores. This finding underscores the efficacy of our approach in expanding the network of thematically related databases far beyond the initial set identified by shared TIDs. It is important to note, however, that while these newly discovered correlations are numerous, they remain a small fraction of all possible pairs. Consequently, their discovery does not substantially bias the negative sampling strategy used during training.
III-D Graph Construction
Upon obtaining the pretrained embedding model, denoted as $f_{\theta^{*}}$, the embedding vector for each database in the corpus can be derived. To construct the database graph, we then compute the cosine similarity between the embedding vectors of all possible database pairs. This computation is performed in parallel for efficiency.
To set the similarity threshold $\tau$ for deciding whether two databases are correlated, we first choose an initial threshold $\tau_{0}$ that maximizes the Youden’s index. To increase confidence in the presence of an edge, we then increase $\tau_{0}$ by $10\%$, $15\%$, and $20\%$ to obtain candidate thresholds $\tau$ for edge determination. Selecting $\tau$ entails a trade-off: a lower value may raise recall for truly related pairs but reduce precision by introducing spurious connections and increase computational costs for subsequent graph analyses; a higher value typically improves precision and confidence but yields a sparser graph. By default, we set $\tau\geq 0.9436$ (i.e., $(1+15\%)\,\tau_{0}$) To achieve a balance between graph density and edge utility.
IV Dataset Details
This section introduces the structural details of the WikiDBGraph dataset (Section IV-A) and subsequently describes the properties defined for the nodes and edges within this graph (Section IV-B). From these properties, we conclude three characteristics of real-world data for CL tasks.
IV-A Graph Structure
The structural characteristics of the generated database graphs, as detailed in Table IV and further visualized by the distributions presented in Figs. 4(a), 4(b), and 4(c), reveal the first characteristic of data in CL tasks.
@@ IV-A Graph Structure
The mean degree reported in Table IV is in the hundreds, indicating extensive interconnection among databases. This is further supported by Fig. 4(a), which shows a substantial share of high-degree databases. However, the network is heterogeneous: the degree distribution is long-tailed—only a small fraction of databases are high-degree, whereas many reside in small clusters or communities, as evidenced by the connected-component size distribution in Fig. 4(b) and the community-size distribution in Fig. 4(c). This imbalance is further visualized in Fig. 3(c), which reveals many small clusters. Such long-tailed structure with a few high-degree “hubs” is common in real-world networks such as social networks [42] and webpage networks [43]. Collectively, these results show that the databases in WikiDBGraph are broadly interconnected, yet connectivity varies across databases, underscoring the necessity of a graph structure for CL studies.
IV-B Node and Edge Properties
To further characterize data in CL tasks, we compute properties for both nodes (databases) and edges (relationships between databases) in the graph, grouped into structural, semantic, and statistical categories. Structural properties capture architectural characteristics (e.g., table counts and schema similarity). Semantic properties encode conceptual meaning (e.g., topic embeddings and their cosine similarity). Statistical properties quantify the underlying data (e.g., column cardinality and distribution divergence). Tables V and VI summarize these properties; specific details follow.
Nodes.
@@ IV-B Node and Edge Properties
Table VI indicates limited feature overlap across databases: the ratios of table and column overlaps are low on average (JaccardTable mean $0.12$; JaccardColumn mean $0.36$), and absolute overlaps remain modest (CommonTables mean $1.85$; CommonColumns mean $74$). Meanwhile, the instance overlap is also limited: the OverlapRatio averages $0.22$, indicating that shared columns contain relatively few identical values across databases. Taken together, these observations imply that most cases cannot be idealized as purely horizontal (same features, disjoint samples) or purely vertical (same samples, disjoint features); rather, the alignment across databases is partial and hybrid.
@@
Although the mean DataVolume per database is small (1.7 MB; Table V), the average AllJoinSize, which is the estimated row count from joining all tables within a database, is $9.5\times 10^{16}$. Such joins are computationally prohibitive in both time and memory, rendering CL methods that assume a single, pre-joined table impractical in this setting. Consequently, practical CL must operate over relational structures directly, avoiding explicit full joins while leveraging graph-aware strategies.
TABLE IV: Summary of Graph Statistical Properties ($C$: Connected Component)
@@
1The ratio of $\tau$ increased by $\tau_{0}$; larger ratios maintains higher confidence of relationship but retain fewer edges. (Default 15%)
TABLE V: Summary of Node (Database) Properties in WikiDBGraph
@@
TABLE VI: Summary of Edge (Database Relationship) Properties in WikiDBGraph ($\tau=0.94$)
@@
V Experiments
This section evaluates the utility of WikiDBGraph for collaborative learning. Section V-A details the experimental settings. Because existing CL algorithms [44, 45] are not directly applicable end-to-end to WikiDBGraph, we design an automated data-mining pipeline (Section V-B) that interfaces with existing CL training algorithms. The performance on the most similar database pairs are presented in Section V-C.
V-A Experimental Settings

### v3

@@ I Introduction
The limitations of existing CL benchmarks are illustrated by a real example from WikiDBs [16], a recently released large-scale database corpus: a collaboration between a National Monuments database and a Historic Places Registry (Fig. 1). Standard benchmarks typically approximate such a setting by perfectly partitioning a single table, implicitly assuming ideal schema and instance alignment. In practice, independently curated web databases exhibit structured yet imperfect overlap: some databases share identifiable attributes such as Latitude and Longitude, but differ in schema (e.g., ArchitecturalStyle vs. EstablishmentDate) and overlap only partially at the instance level—some monuments are historic places, but not all. As detailed later in our hybrid-overlap case study (Section VII-D), this specific component bridges two horizontally aligned clusters, enabling practical collaborative tasks such as predicting national-treasure categories for historic places using these partially vertical connections. This example motivates web-based data for CL evaluation: decentralized curation (e.g., via Wikidata) naturally produces autonomous, weakly aligned databases that more faithfully proxy real-world enterprise silos, thereby exposing essential data-management challenges—such as schema matching and fuzzy joining—that synthetic, perfectly aligned benchmarks [4] typically omit.
@@
By investigating WikiDBs [16], we identify three common yet impractical assumptions in existing CL benchmarks and algorithm design: (1) isolation: each database is treated as an isolated data source, and connections between clients are ignored; (2) alignment: databases are assumed to be aligned either horizontally (disjoint samples, same features) [17] or vertically (disjoint features, same samples) [18]; and (3) joinable: databases can be fully joined into a single database.
@@
We address these challenges by constructing WikiDBGraph, a large-scale, open-source graph of relational databases built from Wikidata [19]. To mitigate the limited explicit correlations, we leverage known relations within WikiDBs [16] and employ contrastive learning to train a model that predicts similarity between databases. The resulting similarity scores induce a graph whose nodes represent databases and whose weighted edges quantify their correlation. To address unaligned and unjoinable databases, we enrich the graph by defining and computing a set of properties for both nodes and edges, derived from database structure, schemas, and data distributions. Using these properties, we propose an automated data-mining pipeline for CL. While this pipeline yields average performance improvements over training on isolated databases, it reveals a critical performance gap compared to ideal centralized training. The contributions of this work include:
@@
    We construct WikiDBGraph, a large-scale graph of 100,000 interconnected relational databases, annotated with 13 node properties and 12 edge properties derived from database structure, to serve as a real-world benchmark for collaborative learning.
@@
    We conduct case studies on feature-, instance-, and hybrid-overlapped database pairs, summarizing key challenges to guide future research.
@@
The paper is organized as follows. Section II introduces background and motivation of WikiDBGraph; Section III surveys database corpora and CL datasets; Section IV details the graph construction method; Section V analyzes graph structure and node/edge properties; Section VI evaluates the benchmark utility and downstream task performance; Section VII presents feature-, instance-, and hybrid-overlap CL case studies. Section VIII concludes the paper. For more details, this technical report also includes the extension to GitTables [22] (Section IX), the performance relation to database size (Section X), findings and open challenges (Section XI), and future directions (Section XII).
II Background
Collaborative learning (CL) trains models across clients without sharing raw data, spanning federated [5, 6, 8, 7, 23, 24], transfer [9], and split learning [10, 25, 26]. Federated learning iteratively aggregates client updates [27], using methods like FedProx [28] and FedOV [29] to mitigate data heterogeneity. When a single global model is insufficient, personalized federated learning [30, 31] tailors models to individual clients. For instance, applying SFL [30] to Fig. 1 allows the National Monuments database to personalize its model by aggregating messages from connected registries over the correlation graph.
Beyond federated learning, split learning [10] achieves collaboration by partitioning a model across clients (and/or a server) and training it through the exchange of intermediate representations and gradients; inference likewise requires the participating parties. This paradigm is particularly well-suited to feature-partitioned settings, where different clients hold complementary feature subsets. Transfer learning [9] provides another collaborative mechanism, often enabling one-shot or low-interaction knowledge transfer: one client trains a model and shares it with another client for adaptation via fine-tuning. While effective in certain communication-constrained scenarios, such one-way transfer typically offers weaker mutual adaptation than multi-round protocols such as federated learning or split learning.
Existing CL systems such as FATE [32] have unified multiple privacy-preserving paradigms—including FL, split learning, and transfer learning—and support diverse models, including linear regression, gradient-boosting decision trees, and support vector machines. Despite this progress, real-world CL deployments remain limited, largely because most algorithms target the model-training stage rather than the end-to-end data management pipeline.
III Related Work
@@
Existing collaborative learning (CL) benchmarks for federated learning (FL) and split learning (SL) often inadequately capture real-world, database-centric collaboration. Many are synthetic, formed by partitioning a centralized dataset to simulate distribution (e.g., FedNoisy [33], FPLLib [34], MarsFL [35], NIIDBench [4], SLPerf [36], VertiBench [13]), and thus may not reflect natural heterogeneity. Others are built from real-world sources (e.g., LEAF [11], ORAF [12], FedML [14], NUS-WIDE [37]) but assume perfect alignment into horizontal CL (disjoint samples, same features) or vertical CL (disjoint features, same samples) [17], whereas practical collaboration often involves partial overlap in both samples and features. Moreover, most benchmarks (i) model each client as a single table and (ii) implicitly treat all inter-client relationships as equally relevant, which is unrealistic for multi-table databases and heterogeneous cross-client correlations. Built from real-world Wikidata, WikiDBGraph is designed to capture these practical complexities by supporting hybrid relationships and varying degrees of correlation across clients.
@@
WikiDBGraph Real Horizontal and Vertical Database
@@
Early tabular corpora primarily consist of isolated tables (e.g., GitTables [22]), which do not capture the relational structure inherent to real-world databases. To better reflect practice, database corpora with multiple interrelated tables have emerged, spanning schema-only resources (e.g., SchemaDB [40], GitSchemas [41]), data-only resources (e.g., SQLShare [39]), and schema-data corpora. However, existing schema-data corpora are often small-scale (e.g., the CTU Prague Relational Learning Repository [43]) or tailored to specific tasks (e.g., Text-to-SQL benchmarks such as Spider [44] and BIRD [42]), limiting their utility for general-purpose representation learning.Despite their scale, recent schema-data corpora such as SchemaPile [45] and WikiDBs [16] largely lack explicit relationships between databases, which hinders studying cross-database collaborative learning. WikiDBGraph addresses this gap by systematically constructing inter-database connections, enabling large-scale CL across distinct databases.
IV Dataset Construction
This section details the construction of WikiDBGraph (Fig. 2). We define the database relationship identification task (Section IV-A), present our approach (Section IV-B), evaluate the embedding model (Section IV-C), and describe the final graph construction (Section IV-D).
IV-A Problem Definition
@@ IV-B Approach
To mitigate the verbosity of the original JSON files and reduce input size for subsequent processing, we serialize each $s_{i}$ and samples in $D_{i}$ into a concise textual format, denoted as abstract $t_{i}\in\mathcal{T}$. This serialization retains both structural information and an abstract of the data content. Specifically, for each database, we preserve its name, the names of its tables, and for each table, the names of its columns along with $\phi$ representative sample values (default $\phi=3$) from each column. An illustrative example of our serialization format is presented below:
@@
While sample values are included to offer a qualitative indication of the data, we do not incorporate full data distributions or the entirety of the raw data. This approach is adopted for two primary reasons: 1) schema-level information (database, table, and column names) combined with data samples is often more directly indicative of the database’s topic than comprehensive statistical distributions of all values. 2) Processing and representing the complete data for all databases would be computationally expensive and lead to excessively large serialized representations. While our manual inspection confirms that cryptic metadata (e.g., column_1) is rare in WikiDBGraph due to upstream curation [16], we acknowledge this remains a challenge in broader settings. In such scenarios, our graph formulation offers a natural mitigation: the semantics of opaque schemas can be inferred from the context of well-defined neighboring databases. We leave the empirical evaluation of this resilience for future work.
@@
We employ a contrastive learning [46] framework to train an embedding model, $f_{\theta}(\cdot):\mathcal{T}\rightarrow\mathbb{R}^{d}$, which maps serialized database schema from the input space $\mathcal{T}$ into a $d$-dimensional vector space. The distinction between positive and negative pairs, crucial for the contrastive loss, is determined by TIDs: a pair of database abstracts $(t_{i},t_{j})$ is a positive pair if they share the same TID; otherwise, it is a negative pair. The sets of positive and negative pairs are partitioned into training, validation, and test subsets in a 7:1:2 ratio, ensuring no database overlap between partitions.
@@
IV-C Evaluation of Embedding Model
We analyze the similarity distributions in Fig. 3(b) to validate the reliability of our learned correlations. The distribution of explicit TID-matched pairs is heavily skewed toward $1.0$, confirming that TIDs serve as high-precision labels with negligible false positives. Conversely, the false negatives are very prevalent according to the global distribution revealing a massive tail of high-similarity pairs lacking TID links. Consequently, leveraging these high-precision (low–false-positive) labels, our model can safely expand beyond TID links to uncover a large set of implicit correlations that TID matching misses, substantially enlarging the database network. Importantly, while these newly discovered correlations are numerous, they remain a small fraction of all possible pairs. Consequently, their discovery does not substantially bias the negative sampling strategy used during training.
@@ IV-C Evaluation of Embedding Model
This learned embedding model is not universal across all database corpora. However, the underlying training paradigm readily extends to other corpora, such as GitTables [22] without pre-defined labels (TIDs). We demonstrate this extension in the technical report [51].
IV-D Graph Construction
Upon obtaining the pretrained embedding model, denoted as $f_{\theta^{*}}$, the embedding vector for each database in the corpus can be derived. To construct the database graph, we perform nearest-neighbor retrieval in the embedding space and compute cosine similarities between each database embedding and its nearest neighbors. We use the faiss library [52] to enable efficient large-scale nearest-neighbor search.
Threshold methodology. We treat $\tau$ as a graph sparsification parameter rather than a tuned hyperparameter. We first obtain a data-driven baseline $\tau_{0}$ from the embedding model’s similarity predictions, using the threshold that maximizes Youden’s index. As shown in Fig. 6, this choice can produce a very dense graph, which may be undesirable for applications that require efficient construction and analysis. Accordingly, we release (i) all graph-agnostic node/edge properties computed over edges with similarity $\geq\tau_{0}$, and (ii) the complete graph-construction pipeline, so that users can instantiate graphs under any threshold suited to their budget and use case. In the paper, we additionally report three representative sparsity regimes (low/medium/high) to illustrate how key graph statistics vary with edge pruning (Table III).
Downstream evaluation and pair selection. Our downstream experiments (Section VI) focus on the highest-similarity pairs because this aligns with how most CL methods are intended to be used: performance is meaningful and comparable only when the paired parties exhibit strong semantic relatedness (e.g., substantial feature/instance overlap). Accordingly, we evaluate on a fixed set of top-ranked pairs ($\tau>0.98$) whose similarities are uniformly very high, making the downstream results effectively insensitive to the particular global sparsification threshold used for graph construction; $\tau$ mainly affects density-related graph analyses and the exploration of more challenging weak-overlap regimes.
V Dataset Details
This section introduces the structural details of the WikiDBGraph dataset (Section V-A) and subsequently describes the properties defined for the nodes and edges within this graph (Section V-B), from which three characteristics emerge.
V-A Graph Structure
The structural characteristics of the generated database graphs, as detailed in Table III and further visualized by the distributions presented in Figs. 4(a), 4(b), and 4(c), reveal the first characteristic of data in CL tasks.
@@ V-A Graph Structure
The mean degree reported in Table III is in the hundreds, indicating extensive interconnection among databases. This is further supported by Fig. 4(a), which shows a substantial share of high-degree databases. However, the network is heterogeneous: the degree distribution is long-tailed—only a small fraction of databases are high-degree, whereas many reside in small clusters or communities, as evidenced by the connected-component size distribution in Fig. 4(b) and the community-size distribution in Fig. 4(c). This imbalance is further visualized in Fig. 3(c), which reveals many small clusters. Such long-tailed structure with a few high-degree “hubs” is common in real-world networks such as social networks [53] and webpage networks [54]. Collectively, these results show that the databases in WikiDBGraph are broadly interconnected, yet connectivity varies across databases, underscoring the necessity of a graph structure for CL studies.
V-B Node and Edge Properties
To further characterize data in CL tasks, we compute properties for both nodes (databases) and edges (relationships between databases) in the graph, grouped into structural, semantic, and statistical categories. Structural properties capture architectural characteristics (e.g., table counts and schema similarity). Semantic properties encode conceptual meaning (e.g., topic embeddings and their cosine similarity). Statistical properties quantify the underlying data (e.g., column cardinality and distribution divergence). Tables IV and V summarize these properties; specific details follow.
Nodes
@@ V-B Node and Edge Properties
Table V indicates limited feature overlap across databases: the ratios of table and column overlaps are low on average (JaccardTable mean $0.12$; JaccardColumn mean $0.36$), and absolute overlaps remain modest (CommonTables mean $1.85$; CommonColumns mean $74$). Meanwhile, the instance overlap is also limited: the OverlapRatio averages $0.22$, indicating that shared columns contain relatively few identical values across databases. Taken together, these observations imply that most cases cannot be idealized as purely horizontal (same features, disjoint samples) or purely vertical (same samples, disjoint features); rather, the alignment across databases is partial and hybrid.
@@
Although the mean DataVolume per database is small (1.5 MB; Table IV), the average AllJoinSize, which is the estimated row count from joining all tables within a database, is $9\times 10^{16}$. Such joins are computationally prohibitive in both time and memory, rendering CL methods that assume a single, pre-joined table impractical in this setting. Consequently, practical CL must operate over relational structures directly, avoiding explicit full joins while leveraging graph-aware strategies.
TABLE III: Summary of Graph Statistical Properties ($C$: Connected Component)
@@
1The ratio of $\tau$ increased by $\tau_{0}$; larger ratios maintains higher confidence of relationship but retain fewer edges.
TABLE IV: Summary of Node (Database) Properties in WikiDBGraph
@@
TABLE V: Summary of Edge (Database Relationship) Properties in WikiDBGraph ($\tau=0.94$)
@@
VI Experiments
This section evaluates the utility of WikiDBGraph for collaborative learning. Section VI-A details the experimental settings. Because existing CL algorithms [55, 56] are not directly applicable end-to-end to WikiDBGraph, we design an automated data-mining pipeline (Section VI-C) that interfaces with existing CL training algorithms. The performance on the most similar database pairs are presented in Section VI-D.
VI-A Experimental Settings